Hot dogs are sold in packages of 12 and hot dog buns are sold in packages of 8. If they predict that they will need at least 150 hot dogs for the event, how many packages of hot dogs and buns should they buy, if every hot dog must be paired with a bun?
Answer by stanbon(75887)

Hot dogs are sold in packages of 12 and hot dog buns are sold in packages of 8. If they predict that they will need at least 150 hot dogs for the event, how many packages of hot dogs and buns should they buy, if every hot dog must be paired with a bun?
-----
# of bun packs = 150/8 = 75/4 = 18.75
Round up to at least 19 bun packs
----
# of hot dog backs = 150/12 = 12.5
Round up to at least 13 hot dog packs
===============
Cheers,
Stan H.
---------------------